OpenStudy (anonymous):

find the center, foci, and vertices of the ellipse. (x-4)^2 +9(y+1)^2=36 don't know how to put it in the the basic form

OpenStudy (phi):

divide both sides of the equation (all terms) by 9

OpenStudy (anonymous):

so it will look like (x-4)^2/9 + (y+1)^2 = 4

OpenStudy (phi):

oops, we want a 1 on the right side, so we should divide both sides by 36 (not 9)

OpenStudy (phi):

that will give a^2= 36 and b^2= 4

OpenStudy (anonymous):

so it will look like (x-4)^2/36 + (x+1)^2/4 =1

OpenStudy (phi):

yes, that is standard form (except you mean (y+1)^2/4 right?) center will be at (4,-1)

OpenStudy (anonymous):

yes!

OpenStudy (phi):

focus is "f" away from the center f= sqrt(a^2-b^2) see http://en.wikipedia.org/wiki/Ellipse#Equations

OpenStudy (anonymous):

so the vertex are (10,-1) and (-2,-1)